Calculate the apparent power in kVA for a 200 amp, 120/208 volt, 3-phase service.

Select one:

a. 769 kVA

b. 72 kVA

c. 576 kVA

d. 432 kVA
Answer #1

In three phase service for line to line voltage , apparent power is given by

S(kVA) = 3 × I(A) × VL-L(V) / 1000

= 3 × 200 × 208 / 1000 = 72.053 kVA

And for line to nutral voltage , apparent power is given by

S(kVA) = 3 × I(A) × VL-N(V) / 1000

=  3 × 200 × 120 / 1000 = 72 kVA

So , the answer is 72 kVA